A second hand textbook store owner wants to increase sales to students by inserting a new disk into texts for courses that require students use additional resources on the disk. This will enable students that purchase the used text to have access to the software license that is normally used up by the original text owner. The store owner has a cost for the disk but is unsure of how long it will take to insert the new disk since this process will be manual and labourious. He decides to conduct some time studies. The books are in boxes, 20 books per box. He observes the worker he has hired to record the times it takes him to replace a box of books. The owner records the time for 10 boxes of books. Since he realizes the worker will get fatigued as the day progresses he builds in a performance rating factor of .93. The worker will also need personal breaks such as going to the bathroom so the owner builds in an allowance factor of .02. T Calculate the standard time of this worker so the owner can know how to calculate a cost for the labour component of this process. The times recorded by the owner for the ten boxes are as follows:

Observations/Time (mins)

1 - 10.23

2 - 10.00

3 - 9.88

4 - 9.97

5 - 10.15

6 - 10.30

7 - 10.24

8 - 11.00

9 - 11.25

10 - 10.88
Answer #1

Mean observed time = (10.23+10+9.88+9.97+10.15+10.3+10.24+11+11.25+10.88)/10 = 10.39 minutes

Normal time = Mean observed time*performance rating factor = 10.39*0.93 minutes = 9.6627 minutes

Standard time = Normal time*allowance factor = 9.6627*1.02 minutes = 9.855954 minutes